OpenStudy (anonymous):

What is the solution to the equation? 4(x - 3) = 7x + 24 A. x = 4 B. x = -36 C. x = -12 D. x = 12

OpenStudy (anonymous):

Distribute 4 into the parenthesis.

myininaya (myininaya):

distribute where you can that is use a(b+c)=ab+ac

OpenStudy (anonymous):

4 x x = 4x. 4 x 3 = ? @zenonisconfuse

OpenStudy (anonymous):

So would the answer be 12?

OpenStudy (anonymous):

That's part of it.

OpenStudy (anonymous):

So that's 4x - 12 = 7x + 24.

OpenStudy (anonymous):

Subtract 24 - 12.

OpenStudy (anonymous):

Okay, now what?

OpenStudy (anonymous):

24 - 12.

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

@zenonisconfuse

OpenStudy (anonymous):

12